(hrw8c2p64) A basketball player, standing near the basket to grab a rebound, jumps 76.6 cm vertically How much time does the player spend in the top 14.8 cm of his jump? Submit Answer Incorrect. Tries 1/5 Previous Tries How much time does the player spend in the bottom 14.8 cm of the jump? Submit Answer Tries 0/5 Does this help explain why such players seem to hang in the air at the top of their jumps? This discussion is closed. Send Feedback
Answer #1

Solution)

Part a)

Here,

t bottom=2*[2*.148/9.8]^0.5 =0.347 second (Ans)

=======

Part b)

t upper =2*[2*.766/9.8]^0.5 -2*[2*.618/9.8]^0.5 =0.084 second

= 0.79- 0.71 = 0.0807 secs (Ans)

======

Yes, it helps as it takes them larger time to come down.